Please clear my confusion about find value of tan7.5°.

Grade:12th pass

3 Answers

Sagar Gupta
7 years ago
What`s your confusion with tan7.5° .Kindly elaborate that where you have stuck?.If you will elaborate then I can help you better.
ANKIT ANAND
7 years ago
Yaa... tan15°=2tan7.5°∕1-tan²7.5° ⇒2-√3=2tan7.5°/1-tan²7.5° . By solving this eq i got value of tan7.5°= (√2-1)(2+√3) . But Ans is (√3-√2)(√2-1). How?
Umakant biswal
7 years ago
@ ankit 
u have to apply the following formula 
sin ( a/2 ) = [(1-cosa)/2] ^ ½ 
cos a/2 = [(1+cosa)/2] ^ ½ 
tan 2a = [2 tan a ] / [1+tan square (a) ] 
answer is = root6-root3 +(root 2 ) -2 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..